OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. Show your work 2 is a factor of n2 - n + 2

OpenStudy (anonymous):

so 2 ^2-2+2 = 4?

OpenStudy (anonymous):

or take comon out

OpenStudy (anonymous):

2n(n+2)?

OpenStudy (perl):

you want to show that 2 divides n^2 -n + 2

OpenStudy (perl):

for any integer n (positive integer)

OpenStudy (perl):

basis case is when n=1 does 2 divide 1^2 - 1 + 2 = why yes, 2 divides 2

OpenStudy (perl):

now you need to show that if its true for some number n=k, then it is true for n=k+1